OpenStudy (anonymous):

Find the absolute value of -3 – 6i.

OpenStudy (anonymous):

l-3-6il i am not sure plz don't mark this answer in your test or whatever u are doing i am not sure.

OpenStudy (anonymous):

uhh ok

OpenStudy (anonymous):

\[a+bi\] has absolute value \(a^2+b^2\)

OpenStudy (anonymous):

oops i mean \[\sqrt{a^2+b^2}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

would i put the negative on the 3?

OpenStudy (anonymous):

in your case it would be \[\sqrt{3^2+6^2}\]

OpenStudy (anonymous):

nope all is positive don't think about minus signs or the \(i\) just use \(\sqrt{a^2+b^2}\)

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

so i'll get sqrt of 45?

OpenStudy (anonymous):

Yeah, sqrt45=sqrt9*sqrt5=3*sqrt5

OpenStudy (anonymous):

i got it, thx anyways